1. If h(x) = 5 for the function h(x) = 2x + 1, what is the value of x?
2
11
12

Found 2 solutions by rfadrogane, Jk22:
Answer by rfadrogane(214)

sol'n:
Just equate the value of h(x)
2x+1 = 5
2x = 5-1
2x = 4, then divide by 2 on both side
so, x = 2---answer

Answer by Jk22(389)

2, since 2*2+1 = 5